OpenStudy (anonymous):

** CALC II ** Find the area of the surface generated by revolving the curve y=sqrt(x+1) from [1,5] about x-axis. I have, as the answer: 2pi[sqrt(6)(sqrt(25/24))] -[sqrt(2)(sqrt(9/8))] but it says it's wrong. I'm not sure where I'm messing up

OpenStudy (anonymous):

\[2\pi[\sqrt(6)(\sqrt(25/24))] -[\sqrt(2)(\sqrt(9/8))]\]

OpenStudy (anonymous):